Owl-Eye Radiologists (OR) does various types of diagnostic imaging. Radiologists perform tests using sophisticated equipment. OR's management wants to compute the costs of performing tests for two different types of patients: those who are hospitalized (including those in emergency rooms) and those who are not hospitalized but are referred by physicians. The data for June for the two categories of patients follow:

Hospital Patients

Other Patients

Total

Units (i.e., procedures) produced

680

880

1,560

Equipment-hours used

230

110

340

Direct labor-hours

500

150

650

Direct labor costs

$

40,000

$

9,000

$

49,000

Overhead costs

$

46,500

The accountant first assigns overhead costs to two pools: overhead related to equipment-hours and overhead related to labor-hours. The analysis of overhead accounts by the cost accountant follows:

Account

Amount

Related to:

Utilities

$

4,790

Equipment-hours

Supplies

12,690

Labor-hours

Indirect labor and supervision

18,510

Labor-hours

Equipment depreciation and maintenance

7,060

Equipment-hours

Miscellaneous

3,450

Equipment-hours

Required:

a. Compute the predetermined overhead rates assuming that Owl-Eye Radiologists uses equipment-hours to allocate equipment-related overhead costs and labor-hours to allocate labor-related overhead costs.

Overhead Rate

Equipment-related

per hour

Labor-related

per hour

b. Compute the total costs of production and the cost per unit for each of the two types of patients undergoing tests in June. (Round "Cost per unit" to 2 decimal places.)

Hospital Patients

Other Patients

Total cost

Cost per unit
Answer #1

a.

Equipment related overheads = Utilities + Equipment depreciation and maintenance + Miscellaneous

= 4,790 + 7,060 + 3,450

= $15,300

Equipment-hours used = 340

Equipment related overhead rate = Equipment related overheads/Equipment-hours used

= 15,300/340

= $45

Labor related overheads = Supplies + Indirect labor and supervision

= 12,690 + 18,510

= 31,200

Direct labor-hours = 650

Labor related overhead rate = Labor related overheads/Direct labor-hours

= 31,200/650

= $48

Overhead rate per hour
Equipment related $45
Labor related $48

b.

Total cost of hospital patients = Direct labor cost + Equipment related overheads + Labor related overheads

= 40,000 + (45 x 230) + (48 x 500)

= 40,000 + 10,350 + 24,000

= $74,350

Cost per unit = Total cost/Number of units produced

= 74,350/680

= $109.34

Total cost of other patients = Direct labor cost + Equipment related overheads + Labor related overheads

= 9,000 + (45 x 110) + (48 x 150)

= 9,000 + 4,950 + 7,200

= $21,150

Cost per unit = Total cost/Number of units produced

= 21,150/880

= $24.03

Hospital patients Other patients
Total cost $74,350 $21,150
Cost per unit $109.34 $24.03